# Merging RLBWTs adaptively

### v2

@@ Abstract
We show how we can merge two run-length compressed Burrows-Wheeler Transforms (RLBWTs) into a run-length compressed extended Burrows-Wheeler Transform (eBWT) in $O(r)$ space and $\tilde{O}(r+L)$ time, where $r$ is the number of runs in the final eBWT and $L$ is the sum of the longest common prefix (LCP) values at the beginnings of those runs.
@@ 1 Introduction
As far as we know, Oliva et al. [7] were the first to study merging RLBWTs, building on Bannai, Gagie and I’s [1] work on dynamic r-indexes. Their merging algorithm uses compressed space but time at least linear in the size of the uncompressed inputs. Recently, Díaz-Domínguez et al. [4] gave an algorithm for merging BWTs that could be extended to RLBWTs and works fairly well in practice but has no good worst-case bound. In this paper we show how we can merge two RLBWTs into a run-length compressed eBWT in $O(r)$ space and $\tilde{O}(r+L)$ time, where $r$ is the number of runs in the final eBWT and $L$ is the sum of the longest common prefix (LCP) values at the beginnings of those runs.
[…]
@@ 2 Preliminaries
For the sake of brevity, we assume readers are familiar with run-length compression, the Burrows-Wheeler Transform (BWT), run-length compressed BWTs (RLBWTs), suffix arrays (SAs) and longest common prefix (LCP) arrays. The extended BWT (eBWT) of two strings $S[1..m]$ and $T[1..n]$ is an interleaving $\mathrm{BWT}_{S,T}[1..m+n]$ of the characters in $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..m]$. For $1\leq h\leq m+n$, $\mathrm{BWT}_{S,T}[h]$ has the lexicographically $h$th largest context. Since $\mathrm{BWT}_{S}$ and $\mathrm{BWT}_{T}$ are subsequences of $\mathrm{BWT}_{S,T}$, if it has $r$ runs then they each have at most $r$ runs.
The context of $\mathrm{BWT}_{S}[i]$ is
@@
(unless $\mathrm{SA}_{T}[j]=1$, in which case the context is just $T[1..n]$), where $\circ$ denotes concatenation. If we assume $S$ and $T$ are each terminated by an end-of-string symbol that occurs nowhere else in $S$ and $T$ then two characters in the same BWT cannot have the same context, and a character in $\mathrm{BWT}_{S}$ and a character in $\mathrm{BWT}_{T}$ have the same context if and only if $S=T$ and the characters have the same positions (in which case we can break ties arbitrarily without affecting $\mathrm{BWT}_{S,T}$).
Given the $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}T[1..n]$ run-length compressed and each with at most $r$ runs, in $\tilde{O}(r)$ time we can build two $O(r)$-space data structures with which, given $i$ and $j$, we can compare $\mathrm{context}_{S}(i)$ and $\mathrm{context}_{T}(j)$ character by character in $\mathrm{polylog}(m+n)$ time per character until we find a mismatch. For example, given $\mathrm{BWT}_{S}$ as a sequence of (character, length) pairs, in $O(r\log r)$ time we can
  1. 1.
    add to each run’s pair its starting position in $\mathrm{BWT}_{S}$, to obtain a triple;
  2. 2.
    stably sort the triples lexicographically by character;
  3. 3.
    add to each triple the prefix sum of the lengths in the triples now before it.
The prefix sum stored in a triple is the value of the $\mathrm{LF}$ function for $S$ applied to the starting position stored in that triple, and the starting position is the value of the $\psi$ function applied to the prefix sum. Therefore, given $i$, we can
  1. 1.
    use binary search to find $i$’s predecessor $\mathrm{pred}(i)$ among the prefix sums in $O(\log r)$ time;
  2. 2.
    take the character stored in the triple with prefix sum $\mathrm{pred}(i)$ as the first character in $\mathrm{context}_{S}(i)$;
  3. 3.
    take the starting position stored in the same triple as $\psi(\mathrm{pred}(i))$;
  4. 4.
    compute the position
    $\psi(i)=\psi(\mathrm{pred}(i))+(i-\mathrm{pred}(i))$
    in $\mathrm{BWT}_{S}$ of the character whose context is $\mathrm{context}_{S}(i)$ cyclically shifted 1 character to the left.
We can repeat this process with $\psi(i)$ to obtain the second character of $\mathrm{context}_{S}(i)$, and so on. We could store a sparse bitvector with 1s marking the prefix sums and use rank and select queries on it to find $\mathrm{pred}(i)$, but we leave such optimizations to the full version of this paper as they change the time complexity by at most polylogarithmic factors.
@@ 3 Algorithm
Suppose we are to merge the RLBWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..n]$ of strings $S[1..m]$ and $T[1..n]$ into the run-length compressed eBWT $\mathrm{BWT}_{S,T}[1..m+n]$ of $S$ and $T$ together. The pseudocode for our algorithm is shown in Figure 1.
@@ 4 Analysis
Since $\mathrm{BWT}_{S}$ and $\mathrm{BWT}_{T}$ both consist of at most $r$ runs, outputting $\mathrm{BWT}_{S}[i..m]$ (line 45) or $\mathrm{BWT}_{T}[j..n]$ (line 47) in run-length compressed form takes $O(r)$ time, where $r$ is the number of runs in $\mathrm{BWT}_{S,T}$, and we can focus on the complexity of the while loop.
It is not difficult to check that during each pass through the while loop, we output an entire run in $\mathrm{BWT}_{S,T}$. Whenever we reach line 9 we have $\mathrm{context}_{S}(i+k)\prec\mathrm{context}_{T}(j^{\prime})$ so, after we output
@@
the next character we output is $\mathrm{BWT}_{S}[i]\neq\mathrm{BWT}_{T}[j]$ and it starts a new run in $\mathrm{BWT}_{S,T}$. It follows that we make $O(r)$ passes through the while loop.
Since $\mathrm{BWT}_{S}$ and $\mathrm{BWT}_{T}$ and $\mathrm{BWT}_{S,T}$ are run-length compressed, the only operations that can take more than constant time during a single pass through the while loop are comparisons between characters’ contexts, of which we make $O(\log(m+n))$ during each pass. It is not difficult to check, however, that whenever we compare two characters’ contexts
@@
$\mathrm{BWT}_{S}[i+k]\neq\mathrm{BWT}_{S}[i]=\mathrm{BWT}_{T}[j]=\mathrm{BWT}_{T}[j+\ell-1]\,.$
The comparison in line 12 is symmetric. When we perform the binary search in line 8, in each step we check whether $\mathrm{context}_{T}(j^{\prime})\succ\mathrm{context}_{S}(i+k)$ for some $j^{\prime}<j+\ell$ with
$\mathrm{BWT}_{T}[j^{\prime}]=\mathrm{BWT}_{T}[j]=\mathrm{BWT}_{S}[i]\neq\mathrm{BWT}_{S}[i+k]\,.$
The binary search in line 13 is symmetric. When we check whether $\mathrm{context}_{S}(i)\prec\mathrm{context}_{T}(j)$ in line 23, $\mathrm{BWT}_{S}[i]\neq\mathrm{BWT}_{T}[j]$. When we check whether $\mathrm{context}_{S}(i+k-1)\prec\mathrm{context}_{T}(j)$ in line 24,
$\mathrm{BWT}_{S}[i+k-1]=\mathrm{BWT}_{S}[i]\neq\mathrm{BWT}_{T}[j]\,.$
The comparison in line 33 is symmetric. When we perform the binary search in line 28, in each step we check whether $\mathrm{context}_{S}(i^{\prime})\succ\mathrm{context}_{T}(j)$ for some $i^{\prime}<i+k$ with
$\mathrm{BWT}_{S}[i^{\prime}]=\mathrm{BWT}_{S}[i]\neq\mathrm{BWT}_{T}[j]\,.$
The binary search in line 37 is symmetric.
It follows that we can charge all the context comparisons to runs in $\mathrm{BWT}_{S,T}$ such that
@@
The number of characters we extract to compare the context of a character in a run $\mathrm{BWT}_{S,T}[p..q]$ to the context of a character in another run is at most
@@
(unless $q=m+n$, in which case the number is just $\mathrm{LCP}_{S,T}[p]$), where $\mathrm{LCP}_{S,T}[1]=0$ and $\mathrm{LCP}_{S,T}[h]$ is the length of the longest common prefix of the $h$th and (h1)(h-1)st characters in $\mathrm{BWT}_{S,T}$. Therefore, the total time we spend comparing contexts is
@@
If we set
@@
then this time bound is $\tilde{O}(r+L)$. Notice we need not compute $L$ to run our algorithm.
Theorem 1
@@
Given the run-length compressed eBWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..n]$, we can merge them into the run-length compressed eBWT $\mathrm{BWT}_{S,T}[1..m+n]$ using $\tilde{O}(r+L)$ time and $O(r)$ space, where $r$ is the number of runs in $\mathrm{BWT}_{S,T}$ and $L$ is the sum of the values in $\mathrm{LCP}_{S,T}$ at the beginnings of runs in $\mathrm{BWT}_{S,T}$.

### v3

@@ Abstract
We show how we can merge two run-length compressed Burrows-Wheeler Transforms (RLBWTs) into a run-length compressed extended Burrows-Wheeler Transform (eBWT) in $O(r)$ space and $O((r+L)\log(m+n))$ time, where $m$ and $n$ are the lengths of the uncompressed strings, $r$ is the number of runs in the final eBWT and $L$ is the sum of the longest common prefix (LCP) values at the beginnings of those runs.
@@ 1 Introduction
As far as we know, Oliva et al. [10] were the first to study merging RLBWTs, building on Bannai, Gagie and I’s [1] work on dynamic r-indexes. Their merging algorithm uses compressed space but time at least linear in the size of the uncompressed inputs. Recently, Díaz-Domínguez et al. [5] gave an algorithm for merging BWTs that could be extended to RLBWTs and works fairly well in practice but has no good worst-case bound. In this paper we show how we can merge two RLBWTs into a run-length compressed eBWT in $O(r)$ space and $O((r+L)\log(m+n))$ time, where $m$ and $n$ are the lengths of the uncompressed strings, $r$ is the number of runs in the final eBWT and $L$ is the sum of the longest common prefix (LCP) values at the beginnings of those runs.
We describe here only how to merge the RLBWTs of two strings. Our algorithm can easily be extended to merge many RLBWTs or run-length compressed eBWTs, but we leave that to the full version of the paper. In Section 2 we review some preliminary concepts, including a technical lemma about Nishimoto and Tabei’s [9] move structures for $\Psi$ with a better bound on the construction time, which may be of independent interest. In Section 3 we present our algorithm and in Section 4 we present a preliminary analysis that yields a time bound of $O((r\log r+L)\log(m+n))$. In Section 5 we prove our technical lemma about building move structures for $\Psi$. Finally, in Section 6 we optimize our algorithm slightly to reduce the $r\log r$ to $r$, so the whole time bound becomes $O((r+L)\log(m+n))$.
@@ 2 Preliminaries
For the sake of brevity, we assume readers are familiar with run-length compression, the Burrows-Wheeler Transform (BWT), run-length compressed BWTs (RLBWTs), suffix arrays (SAs), the $\Psi$ function and longest common prefix (LCP) arrays; see Navarro’s [8] text for an introduction. The extended BWT (eBWT) [7] of two strings $S[1..m]$ and $T[1..n]$ is an interleaving $\mathrm{BWT}_{S,T}[1..m+n]$ of the characters in the BWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..m]$ of $S$ and $T$. Since $\mathrm{BWT}_{S}$ and $\mathrm{BWT}_{T}$ are subsequences of $\mathrm{BWT}_{S,T}$, if it has $r$ runs then they each have at most $r$ runs.
For $1\leq h\leq m+n$, $\mathrm{BWT}_{S,T}[h]$ has the lexicographically $h$th largest context. The context of $\mathrm{BWT}_{S}[i]$ is
@@
(unless $\mathrm{SA}_{T}[j]=1$, in which case the context is just $T[1..n]$), where $\circ$ denotes concatenation. If we assume $S$ and $T$ are each terminated by an end-of-string symbol that occurs nowhere else in $S$ and $T$ then two characters in the same BWT cannot have the same context, and a character in $\mathrm{BWT}_{S}$ and a character in $\mathrm{BWT}_{T}$ have the same context if and only if $S=T$ and the characters are in the same positions (in which case we can break ties arbitrarily without affecting $\mathrm{BWT}_{S,T}$).
Given a character’s position in $\mathrm{BWT}_{S}$ or $\mathrm{BWT}_{T}$, we can extract its context character by character in $O(\log r)$ time plus constant time per character, if we have data structures for iteratively evaluating the $\Psi$ functions for $S$ and $T$ in $O(\log r)$ time plus constant time per iteration. Since $\Psi$ is the inverse of the LF function, Nishimoto and Tabei [9] implicitly showed how to build in $O(r)$ space such structures, which they called move structures, but they said their construction takes $O(m+n+r\log r)$ time. Brown, Gagie and Rossi [3] slightly generalized Nishimoto and Tabei’s result but did not give a construction-time bound at all. Following Brown et al.’s presentation, we show in Section 5 that, given the RLBWTs $\mathrm{BWT}_{S}$ and $\mathrm{BWT}_{T}$, we can build these structures in $O(r\log r)$ time.
Lemma 1.
Given the RLBWTs $\mathrm{BWT}_{S}$ and $\mathrm{BWT}_{T}$, in $O(r\log r)$ time we can build $O(r)$-space data structures for iteratively evaluating the $\Psi$ functions for $S$ and $T$ in $O(\log r)$ time plus constant time per iteration.
It follows that we can compare $\mathrm{context}_{S}(i)$ and $\mathrm{context}_{T}(j)$ in $O(\log r)$ time plus time proportional to the length of their longest common prefix.
@@ 3 Algorithm
Suppose we are to merge the RLBWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..n]$ of strings $S[1..m]$ and $T[1..n]$ into the run-length compressed eBWT $\mathrm{BWT}_{S,T}[1..m+n]$ of $S$ and $T$ together, and we have already built the data structures for $\Psi$ for $S$ and $T$ described in Lemma 1 so we can compare characters’ contexts. The pseudocode for our algorithm is shown in Figure 1.
@@
We exit the while loop only after having consumed all of at least one of $\mathrm{BWT}_{S}$ and $\mathrm{BWT}_{T}$. If we have not consumed all of the other, we output the rest of it (lines 44 to 48).
@@ 4 Analysis
Building the structures in Lemma 1 takes $O(r\log r)\subseteq O(r\log(m+n))$ time, where $r$ is the number of runs in $\mathrm{BWT}_{S,T}$. Since $\mathrm{BWT}_{S}$ and $\mathrm{BWT}_{T}$ both consist of at most $r$ runs, outputting $\mathrm{BWT}_{S}[i..m]$ (line 45) or $\mathrm{BWT}_{T}[j..n]$ (line 47) in run-length compressed form takes $O(r)$ time. Therefore, we can focus on the complexity of the while loop.
Lemma 2.
During each pass through he while loop, we output an entire run in $\mathrm{BWT}_{S,T}$.
Proof.
Whenever we reach line 9 we have $\mathrm{context}_{S}(i+k)\prec\mathrm{context}_{T}(j^{\prime})$ so, after we output
@@
the next character we output is $\mathrm{BWT}_{S}[i]\neq\mathrm{BWT}_{T}[j]$ and it starts a new run in $\mathrm{BWT}_{S,T}$. ∎
It follows from Lemma 2 that we make $O(r)$ passes through the while loop. Since $\mathrm{BWT}_{S}$ and $\mathrm{BWT}_{T}$ and $\mathrm{BWT}_{S,T}$ are run-length compressed, the only operations that can take more than constant time during a single pass through the while loop are comparisons between characters’ contexts, of which we make $O(\log(m+n))$ during each pass, so $O(r\log(m+n))$ in total.
Lemma 3.
Whenever we compare two characters’ contexts
@@
$\mathrm{BWT}_{S}[i+k]\neq\mathrm{BWT}_{S}[i]=\mathrm{BWT}_{T}[j]=\mathrm{BWT}_{T}[j+\ell-1]$
and we output either $\mathrm{BWT}_{S}[i+k]$ or $\mathrm{BWT}_{T}[j+\ell]$ during the next pass. The comparison in line 12 is symmetric. When we perform the binary search in line 8, in each step we check whether $\mathrm{context}_{T}(j^{\prime})\succ\mathrm{context}_{S}(i+k)$ for some $j^{\prime}<j+\ell$ with
$\mathrm{BWT}_{T}[j^{\prime}]=\mathrm{BWT}_{T}[j]=\mathrm{BWT}_{S}[i]\neq\mathrm{BWT}_{S}[i+k]$
and we output $\mathrm{BWT}_{S}[i+k]$ during the next pass. The binary search in line 13 is symmetric. When we check whether $\mathrm{context}_{S}(i)\prec\mathrm{context}_{T}(j)$ in line 23, $\mathrm{BWT}_{S}[i]\neq\mathrm{BWT}_{T}[j]$ and we output either $\mathrm{BWT}_{S}[i]$ or $\mathrm{BWT}_{T}[j]$ during the current pass. When we check whether $\mathrm{context}_{S}(i+k-1)\prec\mathrm{context}_{T}(j)$ in line 24,
$\mathrm{BWT}_{S}[i+k-1]=\mathrm{BWT}_{S}[i]\neq\mathrm{BWT}_{T}[j]$
and we output either $\mathrm{BWT}_{S}[i+k-1]$ during the current pass or $\mathrm{BWT}_{T}[j]$ during the next pass. The comparison in line 33 is symmetric. When we perform the binary search in line 28, in each step we check whether $\mathrm{context}_{S}(i^{\prime})\succ\mathrm{context}_{T}(j)$ for some $i^{\prime}<i+k$ with
$\mathrm{BWT}_{S}[i^{\prime}]=\mathrm{BWT}_{S}[i]\neq\mathrm{BWT}_{T}[j]$
and we output $\mathrm{BWT}_{T}[j]$ during the next pass. The binary search in line 37 is symmetric. ∎
It follows from Lemma 3 that we can charge all the context comparisons to runs in $\mathrm{BWT}_{S,T}$ such that
@@
The number of characters we extract to compare the context of a character in a run $\mathrm{BWT}_{S,T}[a..b]$ to the context of a character in another run is at most
@@
(unless $b=m+n$, in which case the number is just $\mathrm{LCP}_{S,T}[a]$), where $\mathrm{LCP}_{S,T}[1]=0$ and, for $1<h\leq m+n$, $\mathrm{LCP}_{S,T}[h]$ is the length of the longest common prefix of the contexts of the $h$th and (h1)(h-1)st characters in $\mathrm{BWT}_{S,T}$. This takes
@@
time. Therefore, the time for the context comparisons charged to $\mathrm{BWT}_{S,T}[a..b]$ is
@@
Summing over the $r$ runs in $\mathrm{BWT}_{S,T}$, the total time for context comparisons is
$O((r\log r+L)\log(m+n))\,$,
where $L$ is the sum of the LCP values at the beginnings of those runs. Combined with our previous observations, this gives us a preliminary result:
Theorem 4.
@@
Given the run-length compressed eBWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..n]$, we can merge them into the run-length compressed eBWT $\mathrm{BWT}_{S,T}[1..m+n]$ using $O(r)$ space and $O((r\log r+L)\log(m+n))$ time, where $m$ and $n$ are the lengths of the uncompressed strings, $r$ is the number of runs in $\mathrm{BWT}_{S,T}$ and $L$ is the sum of the LCP values at the beginnings of those runs.
5 Proof of Lemma 1
If an RLBWT for a string of length $n$ has $r$ runs then its $\Psi$ function is a permutation on $\{1,\ldots,n\}$ with
$|\{1\}\cup\{i\ :\ 1\leq i\leq n,\Psi(i)\neq\Psi(i-1)+1\}|=r\,.$
Given the RLBWT, in $O(r\log r)$ total time we can first build the set
$\{(1,\Psi(1))\}\cup\{(i,\Psi(i))\ :\ 1\leq i\leq n,\Psi(i)\neq\Psi(i-1)+1\}\,;$
then apply Lemma 5 below to obtain a slightly larger superset $P^{\prime}$ with useful properties; and finally build a move structure with which, given $i$ and $i$’s predecessor in $P^{\prime}$, in constant time we can find $\Psi(i)$ and $\Psi(i)$’s predecessor in $P^{\prime}$. We refer readers to Nishimoto and Tabei’s [9] and Brown et al.’s [3] papers for details about constructing a move structure from $P^{\prime}$; the only significant novelty here is the improved bound on the construction time of $P^{\prime}$. Given only $i$, we can find $i$’s predecessor in $P^{\prime}$ in $O(\log r)$ time with binary search on $P^{\prime}$, then use the move structure to iterate $\psi$ and extract the context of the $i$th character in the RLBWT character by character in constant time per character; Lemma 1 follows. If the original string is over a small alphabet then we can improve the construction time in Lemmas 1 and 5, but it will not affect our final time bound for merging RLBWTs.
Lemma 5.
Let $\pi$ be a permutation on $\{1,\ldots,n\}$ and
$P=\{1\}\cup\{i\ :\ 1<i\leq n,\pi(i)\neq\pi(i-1)+1\}\,.$
Given the set $\{(i,\pi(i))\ :\ i\in P\}$ and an integer $d\geq 2$, in $O(|P|\log|P|)$ time we can build a set $\{(i,\pi(i))\ :\ i\in P^{\prime}\}$, where
  • $P\subseteq P^{\prime}\subseteq\{1,\ldots,n\}$,
  • if $a$ and $b$ are consecutive elements in $\{\pi(i)\ :\ i\in P^{\prime}\}\cup\{n+1\}$ then $|[a,b)\cap P^{\prime}|<2d$,
  • $|P^{\prime}|\leq\frac{d|P|}{d-1}$.
Proof.
Let $Q=\{\pi(i)\ :\ i\in P\}\cup\{n+1\}$. We build AVL trees $T_{P}$ and $T_{Q}$ storing the elements of $P$ and $Q$, respectively. For each $i\in P$ we store $\pi(i)$ as satellite data with $i$ in $T_{P}$, and we store $i$ as satellite data with $\pi(i)$ in $T_{Q}$. For each pair of consecutive elements $a$ and $b$ in $Q$, if $|[a,b)\cap P|\geq 2d$ then we store $a$ in a list $L$. This takes $O(|P|\log|P|)$ total time.
We work in rounds and maintain the relationships between $T_{P}$, $T_{Q}$ and $L$. In each round we check if $L$ is empty and, if so, we return each element $i$ in $T_{P}$ together with its satellite data $\pi(i)$; if not, we remove 1 element from $L$ and then insert 1 element into $T_{P}$, 1 element into $T_{Q}$ and up to 2 elements into $L$.
Let $P_{j}$ and $Q_{j}$ be the contents of $T_{P}$ and $T_{Q}$ after $j$ rounds, so $P_{0}=P$ and $Q_{0}=Q$. Suppose we remove $s$ from $L$ during the (j+1)(j+1)st round, meaning $s,e)\cap P_{j}\geq 2d$, where $e$ is the successor of $s$ in $Q_{j}$. We use $T_{P}$ to find the (d+1)(d+1)st largest element $p$ in $s,e)\cap P_{j}$. We use $T_{Q}$ to find $\pi^{-1}(p)$ by finding the predecessor $q$ of $p$ in $Q_{j}$ and adding pqp-q to $\pi^{-1}(q)$, which is stored as satellite data with $q$. We insert $\pi^{-1}(p)$ into $T_{P}$ with $p$ as satellite data, and insert $p$ into $T_{Q}$ with $\pi^{-1}(p)$ as satellite data. Therefore $Q_{j+1}=\{\pi(i)\ :\ i\in P_{j+1}\}\cup\{n+1\}$.
We use $T_{P}$ and $T_{Q}$ to check first whether $|[p,e)|\cap P_{j+1}|\geq 2d$; if so, we add $p$ to $L$. We then use $T_{P}$ and $T_{Q}$ to find the predecessor $s^{\prime}$ of $\pi^{-1}(p)$ in $Q_{j+1}$ and next element $e^{\prime}$ after $s^{\prime}$ in $Q_{j+1}$, and to check whether $|[s^{\prime},e^{\prime})\cap P_{j+1}|\geq 2d$; if so, we add $s^{\prime}$ to $L$. This completes the (j+1)(j+1)st round, which takes $O(\log|P_{j}|)=O(\log|Q_{j}|)=O(\log(|P|+j))$ total time.
[…]
Since the intervals in $E_{j+1}$ are disjoint and each contain at least $d$ elements of $P_{j+1}$, we have $|P_{j+1}|\geq d|E_{j+1}|\geq d(j+1)$. Since $|P_{j+1}|=|P|+j+1$ and $|P_{j+1}|\geq d(j+1)$, we have $|P|+j+1\geq d(j+1)$ and thus $j+1\leq\frac{|P|}{d-1}$ and $|P_{j+1}|=|P|+j+1\leq\frac{d|P|}{d-1}$. It follows that we return the set $\{(i,\pi(i))\ :\ i\in P^{\prime}\}$ after at most $\frac{|P|}{d-1}\in O(|P|)$ rounds and $O(|P|\log|P|)$ total time. ∎
6 Optimization
The $\log r$ in the time bound in Theorem 4 is the time to find the predecessor in a set $P^{\prime}$ with cardinality $O(r)$ of a position $i$ in an RLBWT, using binary search on $P^{\prime}$. It is multiplied by the $O(r)$ passes we make through the while loop and by the $O(\log(m+n))$ context comparisons we may make during each pass if we reach line 8, 13, 28 or 37 and perform a binary search in an RLBWT. If we do not perform such a binary search during a pass then we make only a constant number of context comparisons in that pass, and they contribute only $O(\log r)$ overhead.
To see how to speed up the binary searches in the RLBWTs, consider how we search for the smallest value $j^{\prime}$ with $\mathrm{context}_{T}(j^{\prime})\succ\mathrm{context}_{S}(i+k)$ in line 8, for example, letting $P_{T}^{\prime}$ be the set we build with Lemma 5 as part of the move structure for $\psi$ for $T$. We repeatedly choose candidate values $j^{\prime}$ in $j,j+\ell)$ and, for each one, compare $\mathrm{context}_{T}(j^{\prime})$ with $\mathrm{context}_{S}(i+k)$. If we choose the candidate values $j^{\prime}$ naïvely, then each context comparison starts with an $O(\log r)$-time predecessor search in $P_{T}^{\prime}$. If we choose $j^{\prime}\in P_{T}^{\prime}$ or already knowing the predecessor of $j^{\prime}$ in $P_{T}^{\prime}$, however, then we do not need such a search.
We first use $O(\log r)$-time binary searches in $P_{T}^{\prime}$ to find the successor of $j$ and the predecessor of $j+\ell-1$. We then use binary search to find the smallest value $p\in[j,j+\ell)\cap P_{T}^{\prime}$ with $\mathrm{context}_{T}(p)\succ\mathrm{context}_{S}(i+k)$, if there is one. Whether $p$ exists or not, we can now focus our search on interval of $j,j+\ell)$ in which all the elements have the same predecessor in $P_{T}^{\prime}$. Binary searches on $\mathrm{BWT}_{S}$ can be sped up symmetrically.
We perform two binary searches instead of one but both are in subsets of $j,j+\ell)$, so Lemma 3 still holds and we can charge all the context comparisons as before. Now, however, the time for the context comparisons charged to a run $\mathrm{BWT}_{S,T}[a..b]$ in $\mathrm{BWT}_{S,T}$ is only
[…]
Summing over the $r$ runs in $\mathrm{BWT}_{S,T}$, the total time for context comparisons is
$O(r\log r+(r+L)\log(m+n))=O((r+L)\log(m+n))\,.$
This gives us our main result:
Theorem 6.
Given the run-length compressed eBWTs $\mathrm{BWT}_{S}[1..m]$ and $\mathrm{BWT}_{T}[1..n]$, we can merge them into the run-length compressed eBWT $\mathrm{BWT}_{S,T}[1..m+n]$ using $O(r)$ space and $O((r+L)\log(m+n))$ time, where $m$ and $n$ are the lengths of the uncompressed strings, $r$ is the number of runs in $\mathrm{BWT}_{S,T}$ and $L$ is the sum of the LCP values at the beginnings of those runs.